# What’s the Weight?

### v1

@@ 1 Person
   3 People 2.55 1.95, 3.33 $<$0.001 2.27 1.66, 3.10 $<$0.001
   4 People 2.51 1.88, 3.36 $<$0.001 2.18 1.57, 3.03 $<$0.001
   5+ People 3.90 2.87, 5.29 $<$0.001 3.51 2.49, 4.94 $<$0.001
@@
   High White Collar 0.73 0.36, 1.49 0.4 1.04 0.48, 2.26 $>$0.9
@@
   Marginally Food Secure 1.80 1.36, 2.38 $<$0.001 2.18 1.59, 2.99 $<$0.001
@@ C.2 Supplemental Analysis of Lead Exposure, SES, and Telomere Length
Table C2: Population Average Treatment Effect (PATE) estimates and corresponding 95% confidence intervals across various analytic approaches for the comparison of effect of lead exposure (high versus low exposure, dichotomized at median $\mu$g/dL) on log-transformed telomere length among $n=5,267$ participants in the National Health and Nutrition Examination Survey (NHANES). IPTW: Inverse Probability of Treatment Weighting.
Method PATE Estimate 95% Confidence Interval
Multiple Regression $0.0002$ $-0.0139,0.0142$
IPTW Estimator $-0.0478$ $-0.0800,-0.0156$
IPTW Estimator w/ Race in Propensity Model $-0.0504$ $-0.0809,-0.0199$
Survey-Weighted Multiple Regression $-0.0032$ $-0.0229,0.0165$
IPTW Multiple Regression $-0.0024$ $-0.0157,0.0109$
IPTW Multiple Regression w/ Race in Propensity Model $-0.0029$ $-0.0161,0.0103$
IPTW + Survey-Weighted Multiple Regression $-0.0051$ $-0.0250,0.0149$
IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model $-0.0053$ $-0.0255,0.0149$
Weighted IPTW + Survey-Weighted Multiple Regression $-0.0049$ $-0.0253,0.0155$
Weighted IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model $-0.0052$ $-0.0257,0.0153$
Proposed OM $0.0001$ $-0.0121,0.0123$
Proposed OM w/ Race in Propensity Model $0.0001$ $-0.0121,0.0123$
Proposed IPW1 $-0.0358$ $-0.0469,-0.0248$
Proposed IPW1 w/ Race in Propensity Model $-0.0368$ $-0.0479,-0.0257$
Proposed IPW2 $-0.0328$ $-0.0433,-0.0223$
Proposed IPW2 w/ Race in Propensity Model $-0.0332$ $-0.0440,-0.0224$

### v2

@@ 1 Person
   3 People 2.55 1.95, 3.33 <<0.001 2.27 1.66, 3.10 <<0.001
   4 People 2.51 1.88, 3.36 <<0.001 2.18 1.57, 3.03 <<0.001
   5+ People 3.90 2.87, 5.29 <<0.001 3.51 2.49, 4.94 <<0.001
@@
   High White Collar 0.73 0.36, 1.49 0.4 1.04 0.48, 2.26 >>0.9
@@
   Marginally Food Secure 1.80 1.36, 2.38 <<0.001 2.18 1.59, 2.99 <<0.001
@@
Figure C1: Absolute standardized mean differences and Kolmogorov-Smirnov statistics for the distribution of socioeconomic status factors before and after adjustment by an unweighted propensity score constructed via a generalized linear model on self-reported race. Dark blue circles represent pre-adjustment statistics, while light blue triangles represent post-adjustment differences. The solid vertical line represents a reference value of 0.0, while the dotted vertical line represents a chosen threshold value of 0.1.
Figure C2: Absolute standardized mean differences and Kolmogorov-Smirnov statistics for the distribution of socioeconomic status factors before and after adjustment by a survey-weighted propensity score constructed via a generalized linear model on self-reported race. Dark blue circles represent pre-adjustment statistics, while light blue triangles represent post-adjustment differences. The solid vertical line represents a reference value of 0.0, while the dotted vertical line represents a chosen threshold value of 0.1.
@@ C.2 Supplemental Analysis of Lead Exposure, SES, and Telomere Length
[…]
We repeated our the setup of our main analysis, treating the dichotomized lead variable as our exposure $(A)$. As before, we fit a propensity score for lead exposure conditional on the twelve socioeconomic indicators $(\boldsymbol{X})$ described in the main text. We further included self-reported race using two strategies – first as a precision covariate along with age and blood composition (i.e., not in $\boldsymbol{X}$) and second, in a sensitivity analysis, as a part of our confounders, $\boldsymbol{X}$. The results of this analysis are presented in Supplemental Table C3. While the point estimates and findings differed across methods, we found a significant exposure effect for both the traditional and proposed IPW estimators, with an estimated 3.3-5.0% reduction in telomere length. For the remaining methods, including the proposed outcome modeling and AIPW estimators, we failed to reject the null hypothesis. Further, there were no appreciable differences in these estimates or the corresponding conclusions regarding significance when comparing our results with versus without including self-reported race in the propensity model (where applicable).
[…]
Method PATE 95% CI
Multiple Regression 0.0002 -0.0139 0.0142
IPTW Estimator -0.0478 -0.0800 -0.0156
IPTW Estimator w/ Race in Propensity Model -0.0504 -0.0809 -0.0199
Survey-Weighted Multiple Regression -0.0032 -0.0229 0.0165
IPTW Multiple Regression -0.0024 -0.0157 0.0109
IPTW Multiple Regression w/ Race in Propensity Model -0.0029 -0.0161 0.0103
IPTW + Survey-Weighted Multiple Regression -0.0051 -0.0250 0.0149
IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model -0.0053 -0.0255 0.0149
Weighted IPTW + Survey-Weighted Multiple Regression -0.0049 -0.0253 0.0155
Weighted IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model -0.0052 -0.0257 0.0153
Outcome Modeling and Direct Standardization 0.0001 -0.0121 0.0123
Outcome Modeling and Direct Standardization w/ Race in Propensity Model 0.0001 -0.0121 0.0123
Inverse Probability Weighting 1 -0.0358 -0.0469 -0.0248
Inverse Probability Weighting 1 w/ Race in Propensity Model -0.0368 -0.0479 -0.0257
Inverse Probability Weighting 2 -0.0328 -0.0433 -0.0223
Inverse Probability Weighting 2 w/ Race in Propensity Model -0.0332 -0.0440 -0.0224
Augmented Inverse Probability Weighting -0.0007 -0.0121 0.0106
Augmented Inverse Probability Weighting w/ Race in Propensity Model -0.0009 -0.0124 0.0107